A coin tossed 900 times . Calculate the standard normal variate for the number of heads between 435 and 465

1
Expert's answer
2022-01-11T12:50:08-0500

"n=900"

"p=0.5"

"\mu =np=900\times0.5=450"


"\sigma=\sqrt{np(1-p)}=\sqrt{450(1-0.5)}=15"

"P(435<X<465)=\Phi(\frac{465-450}{15})-\Phi(\frac{435-450}{15})"


"P(435<X<465)=\Phi(1)-\Phi(-1)"


"P(435<X<465)=\Phi(1)-(1-\Phi(1))"


"P(435<X<465)=\Phi(1)-1+\Phi(1)"


"P(435<X<465)=2\Phi(1)-1"


"P(435<X<465)=2(0.8413)-1"

"P(435<X<465)=1.6826-1"


"P(435<X<465)=0.6826"